OpenStudy (toxicsugar22):

What rate of interstis required so that $1000 will yield $1900 after 5 years if the interset rate is compounded monthly. And I got 12.9% Now How long will it take a sum of money to triple if it is invested at 10% compounded daily

OpenStudy (heisenberg):

there is a formula for this. do you know it / did you use it for the first question?

OpenStudy (toxicsugar22):

yes it is A=P(1+r/n)^nt

OpenStudy (heisenberg):

ok so: r = .10 n = 365 A = 3*P \[3*P = P*(1+\frac{.1}{365})^{.1*t}\]solve for t.

OpenStudy (toxicsugar22):

so what do we do know

OpenStudy (toxicsugar22):

do i divide by 3

OpenStudy (heisenberg):

well you can start by dividing by P to get rid of it. then you will need to solve for t, which will involve using logarithms.

OpenStudy (toxicsugar22):

so 3=p(1+10/365)^365/t

OpenStudy (toxicsugar22):

is it right

OpenStudy (heisenberg):

no. after dividing by P:\[3= (1+\frac{.1}{365})^{.1*t}\]

OpenStudy (heisenberg):

then you must use logarithms. I use ln which is log_e, the natural log. \[\ln(3)= .1*t*\ln(1+\frac{.1}{365})\] then solve for t.

OpenStudy (toxicsugar22):

ok can you help me do that it really hard for me

OpenStudy (toxicsugar22):

please

OpenStudy (heisenberg):

so remember, what you do to one side of the equation, you must do to the others. since we want t by itself, we can divide by the other terms to isolate it, like so: \[\frac{\ln(3)}{\ln(1+\frac{.1}{365})} = .1*t\] then divide by .1 and you will have t by itself.

OpenStudy (toxicsugar22):

ok can u help more

OpenStudy (heisenberg):

what do you not understand?

OpenStudy (toxicsugar22):

the log

OpenStudy (toxicsugar22):

its so complicated

OpenStudy (heisenberg):

can you use a calculator?

OpenStudy (toxicsugar22):

maybe if you go step by step with me then i will get it

OpenStudy (toxicsugar22):

yes i can use a calc.

OpenStudy (toxicsugar22):

i can but i still need help

OpenStudy (heisenberg):

take the log(3), then divide by the other log. then divide by .1

OpenStudy (toxicsugar22):

ok log 3 is .47712125

OpenStudy (toxicsugar22):

now i divide by what

OpenStudy (heisenberg):

log(1+.1/365)

OpenStudy (toxicsugar22):

i got 1.18968

OpenStudy (toxicsugar22):

is it right

OpenStudy (toxicsugar22):

is it right

OpenStudy (toxicsugar22):

i got .4010500723

OpenStudy (heisenberg):

how did you get 1.18968? you must do the inside of the parentheses first. here is it in an online calculator: http://www.wolframalpha.com/input/?i=log%281%2B.1+%2F+365%29&a=*FunClash.log-_*Log10.Log-

OpenStudy (toxicsugar22):

ok But isnt it .10 instead of .1

OpenStudy (toxicsugar22):

because it wat to know How long will it take a sum of money to triple if it is invested at 10% compounded daily

OpenStudy (heisenberg):

those are the same number! .1 = .10 = .100 = .10000000

OpenStudy (toxicsugar22):

so what is the answere

OpenStudy (heisenberg):

can you not calculate it? it would be better then me just giving it to you.

OpenStudy (toxicsugar22):

i tried

OpenStudy (toxicsugar22):

ut i got a weird answer and im not sure i got it right

OpenStudy (toxicsugar22):

it 40104.8

OpenStudy (toxicsugar22):

is that it

OpenStudy (heisenberg):

that doesn't seem right though, heh. does it?

OpenStudy (toxicsugar22):

no it doesnt

OpenStudy (heisenberg):

because i messed up! it is n*t, not r*t. so you should not divide by .1, but by 365

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

so now can u help again

OpenStudy (heisenberg):

you should try and do these yourself on a calculator. you can type it in on this site if you want even. http://www.wolframalpha.com/input/?i=log%283%29+%2F+log%281%2B.1+%2F+365%29+%2F+365&a=*FunClash.log-_*Log10.Log-

OpenStudy (toxicsugar22):

yes

OpenStudy (toxicsugar22):

You do it also and then we will compare aur answers

OpenStudy (toxicsugar22):

okay

OpenStudy (heisenberg):

i just did it. click on that link. it is an online calculator.

OpenStudy (toxicsugar22):

ok i got 10991.6

OpenStudy (toxicsugar22):

is that what you got

OpenStudy (toxicsugar22):

ok thanks

OpenStudy (toxicsugar22):

im going to give you a medal and be your fan if you be my fan

OpenStudy (toxicsugar22):

do u agree

OpenStudy (toxicsugar22):

u there